OpenStudy (anonymous):

the ratio of men to women working for a company 5 to 8. if there are 104 women working for the company,what is the total number of employees?

OpenStudy (anonymous):

520?

OpenStudy (anonymous):

I got 65 by setting it up as a proportion

OpenStudy (anonymous):

5 : 8 x : 104 x = 65 104 + 65 = 169 total number of employees is 169

OpenStudy (anonymous):

damn dif answers =/ i did 13+104=117 TOTAL EMPLOYEES. 5/8=65/104 4*104=8*65 520=520